what is the equation of the line passing through the points (1,3) and (-3,8)
yours sincerely luke

Answer by mananth(16946)

x1 y1 x2 y2
1 3 -3 8

slope m =(y2-y1)/(x2-x1)
m=-1.25

Plug value of the slope and point (1,3) in
Y =mx+b
3.00=-1.25+b
b= 3 +1.25
b= 4.25
So the equation will be
Y =-1.25x +4.25